### v1

# Adaptive Audio-Visual Speech Recognition via Matryoshka-Based Multimodal LLMs

Abstract
Audio-Visual Speech Recognition (AVSR) leverages both audio and visual modalities to enhance speech recognition robustness, particularly in noisy environments. Recent advancements in Large Language Models (LLMs) have demonstrated their effectiveness in speech recognition, including AVSR. However, due to the significant length of speech representations, direct integration with LLMs imposes substantial computational costs. Prior approaches address this by compressing speech representations before feeding them into LLMs. However, higher compression ratios often lead to performance degradation, necessitating a trade-off between computational efficiency and recognition accuracy. To address this challenge, we propose Llama-MTSK, the first Matryoshka-based Multimodal LLM for AVSR, which enables flexible adaptation of the audio-visual token allocation based on specific computational constraints while preserving high performance. Our approach, inspired by Matryoshka Representation Learning, encodes audio-visual representations at multiple granularities within a single model, eliminating the need to train separate models for different compression levels. Moreover, to efficiently fine-tune the LLM, we introduce three LoRA-based Matryoshka strategies using global and scale-specific LoRA modules. Extensive evaluations on the two largest AVSR datasets demonstrate that Llama-MTSK achieves state-of-the-art results, matching or surpassing models trained independently at fixed compression levels.
Adaptive Audio-Visual Speech Recognition
via Matryoshka-Based Multimodal LLMs
1 Introduction
Audio-Visual Speech Recognition (AVSR) aims to improve the robustness of speech recognition systems by utilizing both audio and visual signals to recognize human speech. The correlation between audio and lip movements enables the model to focus on relevant speech content while discarding ambient or background noise. With the rising demand for robust speech recognition systems and the widespread availability of cameras (e.g., smartphones), numerous studies have explored advancements in AVSR technology. They have investigated different neural architectures (Dupont and Luettin, 2000; Noda et al., 2015; Afouras et al., 2018a; Petridis et al., 2018; Ma et al., 2021; Hong et al., 2022), training methods (Ma et al., 2023a; Hong et al., 2023), and methods using self-supervised pretraining (Shi et al., 2022; Haliassos et al., 2023, 2024b; Hsu and Shi, 2022; Haliassos et al., 2024a).
Collection and processing of the LRS2 and LRS3 datasets, and running of the Whisper (OpenAI) model was done by the Imperial College London authors on Imperial College London systems.
Recently, with the growing popularity and versatility of Large Language Models (LLMs), new efforts have emerged to connect LLMs with speech modeling (Lakhotia et al., 2021; Huang et al., 2024; Park et al., 2024). Specifically, in Auditory Speech Recognition (ASR) and Visual Speech Recognition (VSR), researchers have demonstrated the possibility and effectiveness of LLMs in speech recognition (Chen et al., 2024; Hu et al., 2024b; Ma et al., 2024; Yu et al., 2024; Fathullah et al., 2024; Fang et al., 2024a; Lu et al., 2025; Tan et al., 2024; Yeo et al., 2024). By employing multi-modal speech information, a recent work proposes to adapt LLMs in AVSR as well (Llama-AVSR), attaining state-of-the-art recognition performances (Cappellazzo et al., 2025). A common focus of prior works is reducing the sequence length of speech representations before feeding them into the LLM. Since LLMs have a large number of parameters and speech sequences are much longer than text, directly using speech representations imposes a significant computational burden. At the same time, (Cappellazzo et al., 2025) demonstrate that there is a trade-off between how much we compress the audio-visual speech representations and performance: while higher compression ratios enhance computational efficiency, they lead to a degradation in performance. Therefore, a possible solution is training and distributing different models with compression ratios tailored to individual users’ computational resources.
However, retraining existing models for different compression ratios, each requiring a distinct coarse-to-fine granularity, is time-consuming and impractical. For this reason, we propose to leverage the concept of Matryoshka Representation Learning (MRL) (Kusupati et al., 2022; Kudugunta et al., 2024; Nair et al., 2025) to encode audio-visual information at different granularities using a single model. This concept was recently explored in visual-linguistic understanding and reasoning tasks in (Cai et al., 2024; Hu et al., 2024a), demonstrating that Matryoshka-based large vision-language models can support multi-granular visual processing at inference while achieving performance parity with independently trained models for each compression rate.
For our audio-visual setting, with the aspiration to flexibly decide between computational efficiency and performance at inference time within the same model, we propose Llama-Matryoshka (abbreviated as Llama-MTSK in the rest of the paper), a Matryoshka-based Multimodal LLM which caters to different demands based on specific requirements by training simultaneously audio-visual representations of different granularities. Llama-MTSK first produces audio and video tokens using pre-trained encoders, then reduces their length using average pooling or stacking compression methods at multiple compression rates. Then, unlike the previous works using MRL that directly fine-tune all the LLM’s parameters (Cai et al., 2024; Hu et al., 2024a), we propose three LoRA-based Matryoshka approaches (LoRA [Uncaptioned image] ) to parameter-efficiently fine-tune the LLM (i.e., Llama (Dubey et al., 2024)), which is responsible to generate the transcriptions given the audio-visual tokens and textual prompt. These approaches either employ a single global LoRA to learn audio-visual feature tokens at multiple scales (Multi-Scale LoRA [Uncaptioned image] ), or define multiple LoRAs, each of them focusing on scale-specific audio-visual information (Scale-Specific LoRA [Uncaptioned image] ), or a combination of both (Multi-Scale-Specific LoRA [Uncaptioned image] ). At inference, only the projector and LoRA modules associated with the desired compression rate are activated, ensuring both flexibility and efficiency. Our comprehensive experiments on the two largest AVSR datasets demonstrate that our three proposed methods achieve comparable or better performance than training separate models for each combination of audio-video compression rates. Overall, Llama-MTSK exhibits strong performance results, elastic inference, and computational efficiency under a single set of weights.
Our key contributions are as follows:
  • We propose Llama-MTSK, the first Matryoshka-based Multimodal LLM designed for audio-visual speech recognition. By processing audio-visual tokens with multiple compression levels and granularities, and introducing three Matryoshka-based LoRA modules to efficiently fine-tune the pre-trained LLM, Llama-MTSK is able to dynamically adjust the number of tokens processed during inference using a single model, adapting to varying computational resources or desired accuracy levels.
  • Llama-MTSK achieves state-of-the-art results on LRS2 and LRS3, the two largest AVSR datasets, consistently exceeding the performance of models independently trained at specific compression levels. This trend is observed for the ASR, VSR, and AVSR tasks, across both the evaluated compression techniques and granularities.
Figure 1: Training and inference stages for Llama-MTSK. (Left) During training, we produce audio-visual tokens via pre-trained encoders, followed by specific-scale compression and projection modules. Then, we feed the concatenated audio-visual tokens at multiple scales to the pre-trained Llama-based LLM, which is adapted through one of the three proposed LoRA Refer to caption approaches following the Matryoshka Representation Learning principle. (Right) At inference, Llama-MTSK allows us to change on-the-fly the audio-visual compression rates for each input data conditioned on our specific requirements using the same model architecture and weights, enabling high flexibility. Furthermore, only one projector and one LoRA module are activated at inference (in this figure, those associated with the audio and video compression rates equal to $3$), guaranteeing model’s scalability in training and no extra cost in inference. Refer to caption and Refer to caption represent whether the parameters are trained or kept frozen, respectively.
2 Llama-MTSK
The objective of Llama-MTSK is to train an LLM (Llama-based in our setting) that captures audio and visual information at multiple scales, from coarse to fine, thus providing control over the audio-visual granularity during inference. Consequently, a single “universal” model allows us to dynamically adjust the performance-efficiency trade-off at inference time, according to specific needs (Cai et al., 2024; Hu et al., 2024a).
Llama-MTSK follows the structure of Llama-AVSR (Cappellazzo et al., 2025), the first Multimodal LLM (MLLM) tailored for audio-visual speech recognition, with ad-hoc modifications to support MRL (Kusupati et al., 2022). Llama-MTSK computes audio and video tokens via modality-specific pre-trained encoders, and then input them as prefix tokens to the LLM (together with the textual tokens). This approach, denoted as decoder-only, is adopted by several architectures due to its versatility and flexibility (Liu et al., 2023; Lin et al., 2024; Fang et al., 2024b; Fan et al., 2024; Zong et al., 2024; Zhang et al., 2025b; Lee et al., 2024; Fini et al., 2024; Li et al., 2024; Tong et al., 2024; Yao et al., 2024).
Llama-MTSK consists of three main components: 1) pre-trained audio and video encoders, 2) audio and video compression and projection modules, and 3) an LLM which is parameter-efficiently fine-tuned via ad-hoc LoRA-based strategies (i.e., LoRA [Uncaptioned image] ).
2.1 Audio/Video Pre-Trained Encoders
We use pre-trained audio and video encoders to project the input audio and video data into two sets of audio and video tokens. We denote with $\mathbf{X}^{\mathsf{A}}\in\mathbb{R}^{N_{\mathsf{A}}\times d_{\mathsf{A}}}$ and $\mathbf{X}^{\mathsf{V}}\in\mathbb{R}^{N_{\mathsf{V}}\times d_{\mathsf{V}}}$ the audio and video token sequences, respectively, where $N_{\mathsf{A}}$/$N_{\mathsf{V}}$ is the number of audio/video tokens, and $d_{\mathsf{A}}$/$d_{\mathsf{V}}$ is the audio/video token dimension. The pre-trained encoders are maintained frozen during the training stage ( [Uncaptioned image] in Figure 1).
2.2 Audio-Visual Compression and Projection
Since the dimensions of audio and video tokens often differ from that of the textual tokens, MLLMs include a projection layer that maps audio and video tokens into the LLM embedding space. It is common to employ either linear projectors (Liu et al., 2023; Luo et al., 2024; Yao et al., 2024; Li et al., 2024; Liu et al., 2024b; Zhang et al., 2025a) or abstractors (e.g., Q-Former, resampler) (Zhu et al., 2023; Li et al., 2023; Cha et al., 2024). In our setting, following (Cappellazzo et al., 2025), we use a two-layer MLP projector.
In addition to this, since the LLM predominantly accounts for the entire computation and memory consumption of the MLLM, it is customary to compress the number of multimodal tokens (in our case audio-visual tokens) by a specific factor in order to find the optimal balance in terms of efficiency and accuracy. For example, (Cappellazzo et al., 2025; Fang et al., 2024a; Ma et al., 2024; Fathullah et al., 2024) stack multiple consecutive tokens along the token hidden dimension to reduce the number of tokens, whereas other methods rely on the Q-Former architecture (Li et al., 2023) using a fixed number of query tokens (Tang et al., 2023; Yu et al., 2024; Zhang et al., 2025b; Cha et al., 2024). However, all these methods need to decide the compression rate to apply beforehand, which means they generate outputs of a single, predetermined length, lacking the ability to modulate the final sequence length. This constraint limits the ability to balance information density and computational efficiency, particularly in resource-constrained deployment scenarios. Alternatively, one could train a separate model for each desired compression rate, but this approach can be time-consuming and cumbersome in practice.
Figure 2: Our three proposed LoRA Matryoshka (LoRA Refer to caption ) approaches. Multi-Scale (MS) LoRA Refer to caption uses a shared global LoRA module for all the audio-visual token scales (in this specific example there are three scales) to fine-tune the pre-trained matrices of the LLM. The Specific-Scale (SS) variant defines a LoRA module tailored to each scale, learning and specializing to a specific scale. The third approach, Multi-Specific-Scale (MSS), combines MS and SS to support both global and specific-scale LoRAs. The global LoRA is responsible to capture relationships that can be shared among different-scale tokens, while specific-scale LoRAs learn tokens based on the specific scale.
In contrast, we propose to compress the audio and video tokens using multiple compression rates, leading to token sequences at multiple scales, and thus different granularities. We explore two different compression methods to reduce the token sequence length: 1) average pooling, and 2) hidden size stacking, where multiple consecutive frames are stacked along the token hidden dimension. Therefore, we decide beforehand a range of G audio compression rates {$a_{1},a_{2},\cdots,a_{\texttt{G}}$} and $T$ video compression rates {$v_{1},v_{2},\cdots,v_{\texttt{T}}$}. We gradually increase the compression rates (i.e., $a_{i+1}>a_{i}$, $i=1,\cdots,\texttt{G}$). With $a_{i}$ we refer both to the compression rate and the corresponding scale interchangeably (e.g., if $a_{i}$ = $4$, then the corresponding sequence would have $\lfloor\frac{N_{\mathsf{A}}}{4}\rfloor$ tokens). We then compress the audio and video tokens using the chosen rates, producing token sequences at multiple scales: [$\mathbf{X}^{\mathsf{A}}_{a_{1}},\mathbf{X}^{\mathsf{A}}_{a_{2}},\cdots,\mathbf{X}^{\mathsf{A}}_{a_{\texttt{G}}}$] and [$\mathbf{X}^{\mathsf{V}}_{v_{1}},\mathbf{X}^{\mathsf{V}}_{v_{2}},\cdots,\mathbf{X}^{\mathsf{V}}_{v_{\texttt{T}}}$].
At this point, each of these sequences are processed by compression rate-specific linear projectors to align the audio-visual and text tokens (see Figure 1).
2.3 LLM Adaptation via LoRA [Uncaptioned image]
The LLM is responsible for generating the corresponding ASR transcription in an auto-regressive fashion given the audio, video, and textual tokens. We define $\mathbf{X}^{\mathsf{AV}}_{ij}$ as the concatenation of audio and video tokens with audio and video compression rates of $a_{i}$ and $v_{j}$, and the prompt textual tokens $\mathbf{X}^{P}$: $\mathbf{X}^{\mathsf{AV}}_{ij}=[\mathbf{X}^{\mathsf{A}}_{a_{i}},\mathbf{X}^{\mathsf{V}}_{v_{j}},\mathbf{X}^{P}]$. To parameter-efficiently align the LLM with the multimodal inputs, we use LoRA modules (Hu et al., 2021) to adapt the query and value projection matrices of each layer. In our setting, the LLM is trained on multiple audio-visual tokens with different scales. We investigate three different strategies to efficiently fine-tune LLM’s pre-trained matrices via LoRA approximation under a MRL setting: 1) Multi-Scale LoRA Matryoshka (MS LoRA [Uncaptioned image] ), 2) Specific-Scale LoRA Matryoshka (SS LoRA [Uncaptioned image] ), and 3) Multi-Specific-Scale LoRA Matryoshka (MSS LoRA [Uncaptioned image] ). These three methods are illustrated in detail in Figure 2.
The MS LoRA [Uncaptioned image] approach uses a single “global” LoRA to approximate the query and value projection matrices of each LLM’s self-attention layer, regardless of the chosen scale and shared by all the input token sequences. For a pre-trained weight matrix $W$, the projection output is computed as follows:
$\mathbf{H}^{\mathsf{AV}}_{ij}\leftarrow\mathbf{X}^{\mathsf{AV}}_{ij}W+s\cdot\mathbf{X}^{\mathsf{AV}}_{ij}W_{\texttt{MS}},$ (1)
where $s$ is a tunable scalar hyperparameter, $W_{MS}=W_{\texttt{MS}}^{down}W_{\texttt{MS}}^{up}$, $W_{\texttt{MS}}^{down}\in\mathbb{R}^{d\times r}$ and $W_{\texttt{MS}}^{up}\in\mathbb{R}^{r\times d}$, and $r\ll d$ ($r$ is the bottleneck dimension).
[…]
$\mathbf{H}^{\mathsf{AV}}_{ij}\leftarrow\mathbf{X}^{\mathsf{AV}}_{ij}W+s\cdot\mathbf{X}^{\mathsf{AV}}_{ij}W_{\texttt{SS}}^{ij},$ (2)
where $W_{\texttt{SS}}^{ij}$ is the LoRA decomposition matrix defined for the $i$-th audio scale and $j$-th video scale, and it is defined as $W_{MS}$. As we explain in subsection 2.4, while all the LoRA modules are used during the training stage, at inference we only activate one LoRA module, corresponding to the selected audio and video scales.
The third approach, MSS LoRA [Uncaptioned image] , is a hybrid approach between MS and SS, which aims to learn both scale-specific and multi-scale audio-visual representations. Consequently, we define both a multi-scale global LoRA module, which is always activated and shared among all the input sequences both at training and at inference, and multiple scale-specific LoRA modules. In this case, the output takes the following form:
$\mathbf{H}^{\mathsf{AV}}_{ij}\leftarrow\mathbf{X}^{\mathsf{AV}}_{ij}W+s\cdot\mathbf{X}^{\mathsf{AV}}_{ij}W_{\texttt{SS}}^{ij}+s\cdot\mathbf{X}^{\mathsf{AV}}_{ij}W_{\texttt{MS}}.$ (3)
Regardless of the LoRA [Uncaptioned image] fine-tuning approach we employ, Llama-MTSK is trained by averaging the auto-regressive next token prediction loss for each audio-visual scale $ij$ for each input data. The LLM predicts the response $\mathbf{Y}=\{y_{l}\}_{l=1}^{L}$ conditioned on the multimodal input tokens, where $L$ represents the number of tokens of the ground truth transcription to be generated. Accordingly, for each Matryoshka audio-visual representation $\mathbf{X}^{\mathsf{AV}}_{ij}$, the probability of the target $\mathbf{Y}$ is computed by:
$p(\mathbf{Y}|\mathbf{X}^{\mathsf{AV}}_{ij})=\prod_{l=1}^{L}p_{\theta}(y_{l}|\mathbf{X}^{\mathsf{AV}}_{ij},y_{<l}),$ (4)
where $y_{<l}$ is the generated output sequence up to token $l-1$, and $\theta$ is the trainable parameters, which comprises the projection layers and the LoRA [Uncaptioned image] modules according to the LoRA [Uncaptioned image] fine-tuning approach used.
The final objective is the average over all the audio-visual token scales:
$\frac{1}{\texttt{G}\cdot\texttt{T}}\sum_{i=1}^{\texttt{G}}\sum_{j=1}^{\texttt{T}}-\log p(\mathbf{Y}|\mathbf{X}^{\mathsf{AV}}_{ij}).$ (5)
2.4 Llama-MTSK: Training vs Inference
During training, Llama-MTSK learns multiple sets of audio-visual tokens, each progressively incorporating more details as the scale increases. To do so, the LLM processes all the multi-scale audio-visual tokens and concurrently optimize over them using Eq. 5. This means that all the projectors and LoRA [Uncaptioned image] modules are involved. Instead, at inference time, for each input data, we choose a specific audio-visual scale and we activate only the projector and LoRA module associated with it. This is equivalent to one single Llama-AVSR model trained on the specific scale. This principle is similar to the behaviour of Mixture of Experts-based models (Shazeer et al., 2017; Fedus et al., 2022; Zoph et al., 2022; Mustafa et al., 2022; Puigcerver et al., 2023; Cappellazzo et al., 2024a; Jiang et al., 2024; Muennighoff et al., 2024), which at inference time only activate a small subset of the available experts (in our case the “experts” are the projectors and LoRA [Uncaptioned image] modules). Figure 1 depicts a schematic comparison of Llama-MTSK training and inference processes.
3 Experiments and Results
3.1 Implementation Details
Datasets. We train and evaluate Llama-MTSK on LRS2 (Son Chung et al., 2017) and LRS3 (Afouras et al., 2018b), the two largest publicly available datasets for audio-visual speech recognition. LRS2 includes $225$ hours of video clips from BBC programs. LRS3 contains $433$ hours of transcribed English video clips from TED talks.
Pre-Processing. We follow (Ma et al., 2023b; Cappellazzo et al., 2025) for the pre-processing of the datasets. For the video modality, we crop the mouth region of interests (ROIs) through a bounding box of $96$ $96$. Each frame is normalised by subtracting the mean and dividing by the standard deviation of the training set. Audio data only undergo z-normalisation per utterance.
Tasks. The AVSR task is studied for the main results, both for LRS2 and LRS3. We also report the results for the ASR and VSR tasks on LRS3.
Llama-MTSK Details. We use Whisper Small and Medium (Radford et al., 2023) as pre-trained audio encoder, whilst AV-HuBERT Large (Shi et al., 2022) for computing the video tokens. Their weights remain frozen throughout the training phase. The projectors consist of two linear layers with ReLU activation in between. As for the LLM, based on the task and dataset, we experiment with $3$ base pre-trained models of varying size from the Llama 3 family (Dubey et al., 2024): Llama 3.1-8B, Llama 3.2-3B, and Llama 3.2-1B. Each LoRA module used to fine-tune the query and key projection matrices of each LLM’s self-attention layer has a bottleneck dimension $r$ such that the original LLM’s hidden size is reduced of a factor $32$ for Llama 3.2-3B and 3.2-1B, and $64$ for Llama 3.1-8B (e.g., for Llama 3.2-1B, since the hidden size is $2048$, the rank is set to $2048/32=64$). The hyperparameter $s$ is set to $\frac{1}{8}$.
Table 1: Comparison between Llama-AVSR and our proposed Llama [Uncaptioned image] MS, SS, and MSS approaches on LRS2 and LRS3 benchmarks. Llama-AVSR trains $4$ independent models tailored to each configuration of audio-video compression rates.
Method Compression Rates (A,V)
(4,2) (4,5) (16,2) (16,5)
LRS3 Dataset
Llama-AVSR 2.4 2.8 3.3 4.1
[…]
Llama [Uncaptioned image] SS 2.3 2.2 3.3 3.6
Llama [Uncaptioned image] MSS 2.4 2.4 3.2 3.5
LRS2 Dataset
Llama-AVSR 4.1 4.5 5.3 8.1
[…]
Llama [Uncaptioned image] SS 3.4 4.7 4.8 6.4
Llama [Uncaptioned image] MSS 3.6 4.8 6.1 9.0
Audio-Visual Token Compression Rates. We choose the audio and video compression rates to train and evaluate Llama-MTSK carefully, based on the studied tasks. For ASR, we apply compression rates in the range of {$4$, $8$, $12$, $16$, $20$}. For VSR, since the task is more challenging, we can afford smaller rates: {$1$, $2$, $3$, $4$, $5$} (we also include the case in which no compression is applied). For AVSR, we apply audio rates in {$4$, $16$} and video rates in {$2$, $5$}, leading to $4$ audio-visual configurations. To compress the audio and video tokens, either we apply average pooling with kernel size and stride equal to the desired compression rate, or we stack consecutive frames along the hidden dimension according to the rate (we denote this as “stacking”).
Table 2: Comparison between Llama [Uncaptioned image] and multiple SOTA methods on the LRS2 and LRS3 benchmarks. The “Lab. Hrs.” column with values X/Y specifies how many labeled hours have been used in training for LRS2 (X) and LRS3 (Y).
Method Rates Lab. Dataset
(A,V) Hrs. LRS2 LRS3
CM-seq2seq (1,1) 380/433 3.7 2.3
Eff. Conf. (1,1) 818/818 2.3 1.8
auto-avsr (1,1) 3448/1902 1.5 1.0
W-Flamingo (1,1) 1982/433 1.4 1.1
USR (1,1) 1982/1759 1.9 1.1
[…]
Llama [Uncaptioned image] SS (4,2) 223/433 2.4 0.9
Llama [Uncaptioned image] MSS (4,2) 223/433 2.4 1.2
[…]
3.2 AVSR Main Results
Figure 3: WER results for the average pooling (left) and stacking (right) compression methods for the ASR task. We use Whisper Small as audio encoder and Llama 3.2-1B as LLM.
We report the results achieved by Llama-MTSK MS, SS, and MSS on the LRS2 and LRS3 datasets in Table 1. We replace “MTSK” with [Uncaptioned image] in the tables and in the following sections to simplify the notation. For both datasets, we use Whisper Small as audio encoder. For the LLM, we use Llama 3.2-1B for LRS3 and Llama 3.2-3B for LRS2. The smaller size of the LRS2 dataset necessitates the larger LLM to mitigate higher WERs. We apply audio compression rates of $4$ and $16$ and video compression rates of $2$ and $5$, resulting in $4$ different compression configurations. We compare these results with those achieved by training Llama-AVSR independently on the $4$ configurations, leading to $4$ models. During inference, Llama-AVSR employs a separate model trained for each audio-video compression rate. In contrast, our Llama [Uncaptioned image] uses a single pre-trained model, activating the projector and LoRA [Uncaptioned image] modules corresponding to the desired compression rate. On the LRS3 dataset, the three proposed Llama [Uncaptioned image] approaches achieve comparable or superior performance to Llama-AVSR, particularly for the SS and MSS configurations. These two methods use LoRA modules specialized for specific compression rates, which are activated during inference based on specific requirements. On the LRS2 dataset, Llama [Uncaptioned image] SS outperforms all other approaches across all compression rates.
Llama [Uncaptioned image] vs SOTA Methods. In Table 2, we compare Llama [Uncaptioned image] with state-of-the-art (SOTA) methods on LRS2 and LRS3 for the AVSR task. We equip Llama [Uncaptioned image] with Whisper Medium and Llama 3.1-8B. We report results from $5$ recent SOTA AVSR methods: CM-seq2seq (Ma et al., 2021), Efficient Conformer (Burchi and Timofte, 2023), auto-avsr (Ma et al., 2023b), Whisper-Flamingo (Rouditchenko et al., 2024), and USR (Haliassos et al., 2024a). Notably, all these methods do not reduce the token sequence length, whereas Llama-AVSR and Llama [Uncaptioned image] reduce the number of tokens by a factor $4$ for audio and $2$ for video. For LRS3, Llama [Uncaptioned image] achieves SOTA results, with its SS variant surpassing Llama-AVSR, which is trained on those specific compression rates, and outperforming methods like auto-avsr and USR, which use more training hours. For LRS2, Llama [Uncaptioned image] SS and MSS perform comparably to Llama-AVSR, while MS achieves better results. Additionally, our methods perform as well as or better than CM-seq2seq and Efficient Conformer but slightly underperform other SOTA methods. However, Llama [Uncaptioned image] is trained only on the $223$ hours of LRS2, whereas all competing methods utilize at least $1982$ hours. We leave for future work the integration of additional training data to enable a fairer comparison. Finally, more AVSR experiments can be found in the Appendix.
3.3 Additional Results
In this section, we extend our analysis to the tasks of ASR and VSR, where only audio or video tokens are fed to the LLM, respectively. We finally present the computational cost analysis of Llama [Uncaptioned image] .
Table 3: Comparison between Llama [Uncaptioned image] MS and a training-free Llama-AVSR-based approach that reduces the number of tokens via average pooling at inference time for the ASR task on the LRS3 dataset.
Method Compression Rate
2 4 6 8 10
Avg Pooling 4.3 13.5 46.1 89.2 160.0
[…]
Figure 4: WER results for the average pooling (left) and stacking (right) compression methods for the VSR task. We use AVHuBERT Large as video encoder and Llama 3.2-3B as LLM.
ASR Results. For the ASR task, we consider $5$ compression rates in the range {$4$, $8$, $12$, $16$, $20$}. In Figure 3, we report the results on the LRS3 dataset when using average pooling compression (left) and stacking compression (right). With the exception of rate = $20$, all the three Llama [Uncaptioned image] methods outperform separately-trained Llama-AVSR methods. The MSS configuration achieves the best WER performance across all the compression rates, even surpassing or equaling the performance of Llama-AVSR trained at the lowest compression rate of $20$.
VSR Results. Figure 4 shows WER results for the VSR task, similar to the ASR results in Figure 3. The video rates are {$1$, $2$, $3$, $4$, $5$}, lower than the ASR rates due to the greater complexity of VSR. For both average pooling and stack compression, all three Llama [Uncaptioned image] approaches outperform Llama-AVSR, with increasing gains at higher rates. The MS and SS approaches using average pooling achieve WER reductions exceeding $10$ at the highest rates. We attribute this improvement at higher compression rates to the joint training of multi-scale tokens. The performance of the three LoRA [Uncaptioned image] approaches varies slightly depending on the compression method, suggesting that no single approach is superior across all configurations. However, all of them significantly outperform Llama-AVSR.
Table 4: Computational cost analysis of Llama [Uncaptioned image] MS using different compression rates and Llama 3.1-8B.
(A,V) Rates # Tokens TFLOPs
(1,1) 757 11.40
[…]
(4,5) 182 2.74
(16,2) 163 2.46
(16,5) 88 1.33
Llama [Uncaptioned image] vs Avg Pooling at Inference Time. Llama [Uncaptioned image] trains a single model that supports multiple scales at inference time by applying different compression rates. We compare our method with a training-free approach that trains a single Llama-AVSR model without compression and then applies the desired compression rate at inference on-the-fly by average pooling the tokens. In Table 3, we study the ASR setting with audio compression rates in the range {$2$, $4$, $6$, $8$, $10$}. The performance of the average-pooling baseline is severely impacted by a decrease in the number of tokens, while Llama [Uncaptioned image] MS is much more robust. These results demonstrate that Llama [Uncaptioned image] MS can be effectively used with diverse computational resources. Notably, even with limited resources, a compression rate of $8$ incurs minimal performance loss.
Computation Cost Analysis. Table 4 presents the computational benefits of using Llama [Uncaptioned image] . Specifically, we evaluate MS LoRA [Uncaptioned image] with Llama 3.1-8B as LLM and detail the associated inference costs. Without compression, we assume the LLM processes $500$ audio tokens, $250$ video tokens (the resolution of the audio encoder is twice that of the video encoder), and $7$ tokens for the textual prompt, totaling $757$. As shown in the table, our proposed approach yields speedups, reducing TFLOPs by over $8$x when applying compression rates of $16$ and $5$ for audio and video, respectively, thus substantially improving efficiency.
4 Conclusion
This work introduces Llama-MTSK, a versatile audio-visual MLLM capable of elastic inference across multiple tasks and computational resources. Llama-MTSK exploits the concept of matryoshka representation learning to adapt the pre-trained LLM through ad-hoc LoRA [Uncaptioned image] modules, achieving performance comparable to or better than models separately trained on each compression rate while significantly reducing computational costs.
5 Limitations
[…]
Appendix A Appendix
Figure 5: Additional WER results using stacking compression for the ASR task with {$2$, $4$, $6$, $8$, $10$} rates. We use the same configuration as in Figure 3.
A.1 Additional Experiments for ASR
In this section, we report additional results for the ASR task when using compression rates in a different range , specifically {$2$, $4$, $6$, $8$, $10$}. Compared to Figure 3, the increment between two consecutive rates is halved. We argue that it is more useful to use more diverse rates for ASR since we do not observe much deterioration of the WER results when doubling the rate (in Figure 5, the baseline Llama-AVSR achieves similar results when compressing the tokens of a factor $2$, $4$, and $6$). Figure 5 shows that Llama [Uncaptioned image] MS and MSS achieves comparable or better performance than Llama-AVSR. As for the SS approach, it performs slightly worse than Llama-AVSR for the first compression rates, and we believe this is because having a specific LoRA module for multiple rates which do not show WER deterioration leads to overfitting as one global LoRA is sufficient. This argument also explains why for rates $8$ and $10$ the MS variant performs better than the other ones.
Figure 6: Additional results for Llama Refer to caption using stacking compression on the LRS3 dataset.
A.2 AVSR Results with Stacking Compression
We include additional results for AVSR on LRS3 using the stacking compression method in Figure 6. The methods use Whisper Small and Llama 3.2-1B as LLM. Our three proposed Matryoshka approaches performs better than or equally well as Llama-AVSR, especially under conditions of high audio compression, underscoring the effectiveness of our proposed Llama [Uncaptioned image] .
A.3 Full AVSR Results with Whisper Medium and LLama 3.1-8B
In Table 2, we only included for Llama-AVSR and Llama [Uncaptioned image] the results with audio and video compression rates equal to $4$ and $2$, respectively. In Table 5, we also report additional configurations of audio-video compression rates. We use Whisper medium as audio encoder and Llama 3.1-8B as LLM. Once more, our proposed methods perform on par or even better than independently-trained Llama-AVSR models for each compression rates configurations. In particular, we highlight the sizeable gains brought by all the three LoRA [Uncaptioned image] approaches for LRS3 when we apply the highest compression rates configuration (16,5).
Table 5: Comparison between Llama-AVSR and our proposed Llama [Uncaptioned image] MS, SS, and MSS approaches on LRS2 and LRS3 benchmarks. We employ Whisper medium and Llama 3.1-8B. Llama-AVSR trains $4$ independent models tailored to each configuration of audio-video compression rates.
Method Compression Rates (A,V)
(4,2) (4,5) (16,2) (16,5)
LRS3 Dataset
Llama-AVSR 0.9 0.9 1.6 2.1
[…]
Llama [Uncaptioned image] SS 0.9 1.0 1.7 1.8
Llama [Uncaptioned image] MSS 1.2 1.0 1.5 1.6
LRS2 Dataset
Llama-AVSR 2.4 2.2 2.9 3.3
[…]
Llama [Uncaptioned image] SS 2.4 2.1 2.9 2.9
Llama [Uncaptioned image] MSS 2.4 2.5 3.2 3.4

### v2

@@
 Abstract
-Audio-Visual Speech Recognition (AVSR) leverages both audio and visual modalities to enhance speech recognition robustness, particularly in noisy environments.
-Recent advancements in Large Language Models (LLMs) have demonstrated their effectiveness in speech recognition, including AVSR.
-However, due to the significant length of speech representations, direct integration with LLMs imposes substantial computational costs.
-Prior approaches address this by compressing speech representations before feeding them into LLMs.
-However, higher compression ratios often lead to performance degradation, necessitating a trade-off between computational efficiency and recognition accuracy.
-To address this challenge, we propose Llama-MTSK, the first Matryoshka-based Multimodal LLM for AVSR, which enables flexible adaptation of the audio-visual token allocation based on specific computational constraints while preserving high performance.
-Our approach, inspired by Matryoshka Representation Learning, encodes audio-visual representations at multiple granularities within a single model, eliminating the need to train separate models for different compression levels.
-Moreover, to efficiently fine-tune the LLM, we introduce three LoRA-based Matryoshka strategies using global and scale-specific LoRA modules.
-Extensive evaluations on the two largest AVSR datasets demonstrate that Llama-MTSK achieves state-of-the-art results, matching or surpassing models trained independently at fixed compression levels.
-Adaptive Audio-Visual Speech Recognition
-via Matryoshka-Based Multimodal LLMs
-1 Introduction
+Audio-Visual Speech Recognition (AVSR) leverages audio and visual modalities to improve robustness in noisy environments.
+Recent advances in Large Language Models (LLMs) show strong performance in speech recognition, including AVSR.
+However, the long speech representations lead to high computational costs for LLMs.
+Prior methods compress inputs before feeding them to LLMs, but high compression often harms accuracy.
+To address this, we propose Llama-MTSK, the first Matryoshka-based Multimodal LLM for AVSR, which flexibly adapts audio-visual token allocation under varying compute constraints.
+Inspired by Matryoshka Representation Learning, our model encodes representations at multiple granularities with a single architecture, avoiding the need for separate models.
+For efficient fine-tuning, we introduce three LoRA-based strategies using global and scale-specific modules.
+Evaluations on major AVSR datasets show Llama-MTSK matches or outperforms models trained at fixed compression levels.
+Index Terms:
+Audio-Visual Speech Recognition, Matryoshka Representation Learning, Elastic Inference
+I Introduction
 Audio-Visual Speech Recognition (AVSR) aims to improve the robustness of speech recognition systems by utilizing both audio and visual signals to recognize human speech.
@@
 With the rising demand for robust speech recognition systems and the widespread availability of cameras (e.g., smartphones), numerous studies have explored advancements in AVSR technology.
-They have investigated different neural architectures (Dupont and Luettin, 2000; Noda et al., 2015; Afouras et al., 2018a; Petridis et al., 2018; Ma et al., 2021; Hong et al., 2022), training methods (Ma et al., 2023a; Hong et al., 2023), and methods using self-supervised pretraining (Shi et al., 2022; Haliassos et al., 2023, 2024b; Hsu and Shi, 2022; Haliassos et al., 2024a).
-Collection and processing of the LRS2 and LRS3 datasets, and running of the Whisper (OpenAI) model was done by the Imperial College London authors on Imperial College London systems.
-Recently, with the growing popularity and versatility of Large Language Models (LLMs), new efforts have emerged to connect LLMs with speech modeling (Lakhotia et al., 2021; Huang et al., 2024; Park et al., 2024).
-Specifically, in Auditory Speech Recognition (ASR) and Visual Speech Recognition (VSR), researchers have demonstrated the possibility and effectiveness of LLMs in speech recognition (Chen et al., 2024; Hu et al., 2024b; Ma et al., 2024; Yu et al., 2024; Fathullah et al., 2024; Fang et al., 2024a; Lu et al., 2025; Tan et al., 2024; Yeo et al., 2024).
-By employing multi-modal speech information, a recent work proposes to adapt LLMs in AVSR as well (Llama-AVSR), attaining state-of-the-art recognition performances (Cappellazzo et al., 2025).
+They have investigated different neural architectures [1, 2, 3, 4, 5, 6], training methods [7, 8], and methods using self-supervised pretraining [9, 10, 11, 12, 13].
+Recently, with the growing popularity and versatility of Large Language Models (LLMs), new efforts have emerged to connect LLMs with speech modeling [14, 15, 16].
+Specifically, in Auditory Speech Recognition (ASR) and Visual Speech Recognition (VSR), researchers have demonstrated the possibility and effectiveness of LLMs in speech recognition [17, 18, 19, 20, 21, 22, 23, 24, 25].
+By employing multi-modal speech information, recent works propose to adapt LLMs in AVSR as well, attaining state-of-the-art recognition performances [26, 27].
 A common focus of prior works is reducing the sequence length of speech representations before feeding them into the LLM.
 Since LLMs have a large number of parameters and speech sequences are much longer than text, directly using speech representations imposes a significant computational burden.
-At the same time, (Cappellazzo et al., 2025) demonstrate that there is a trade-off between how much we compress the audio-visual speech representations and performance: while higher compression ratios enhance computational efficiency, they lead to a degradation in performance.
+At the same time, [26] demonstrate that there is a trade-off between how much we compress the audio-visual speech representations and performance: while higher compression ratios enhance computational efficiency, they lead to a degradation in performance.
 Therefore, a possible solution is training and distributing different models with compression ratios tailored to individual users’ computational resources.
 However, retraining existing models for different compression ratios, each requiring a distinct coarse-to-fine granularity, is time-consuming and impractical.
-For this reason, we propose to leverage the concept of Matryoshka Representation Learning (MRL) (Kusupati et al., 2022; Kudugunta et al., 2024; Nair et al., 2025) to encode audio-visual information at different granularities using a single model.
-This concept was recently explored in visual-linguistic understanding and reasoning tasks in (Cai et al., 2024; Hu et al., 2024a), demonstrating that Matryoshka-based large vision-language models can support multi-granular visual processing at inference while achieving performance parity with independently trained models for each compression rate.
+For this reason, we propose to leverage the concept of Matryoshka Representation Learning (MRL) [28, 29, 30] to encode audio-visual information at different granularities using a single model.
+This concept was recently explored in visual-linguistic understanding and reasoning tasks in [31, 32], demonstrating that Matryoshka-based large vision-language models can support multi-granular visual processing at inference while achieving performance parity with independently trained models for each compression rate.
 For our audio-visual setting, with the aspiration to flexibly decide between computational efficiency and performance at inference time within the same model, we propose Llama-Matryoshka (abbreviated as Llama-MTSK in the rest of the paper), a Matryoshka-based Multimodal LLM which caters to different demands based on specific requirements by training simultaneously audio-visual representations of different granularities.
 Llama-MTSK first produces audio and video tokens using pre-trained encoders, then reduces their length using average pooling or stacking compression methods at multiple compression rates.
-Then, unlike the previous works using MRL that directly fine-tune all the LLM’s parameters (Cai et al., 2024; Hu et al., 2024a), we propose three LoRA-based Matryoshka approaches (LoRA [Uncaptioned image] ) to parameter-efficiently fine-tune the LLM (i.e., Llama (Dubey et al., 2024)), which is responsible to generate the transcriptions given the audio-visual tokens and textual prompt.
+Then, unlike the previous works using MRL that directly fine-tune all the LLM’s parameters [31, 32], we propose three LoRA-based Matryoshka approaches (LoRA [Uncaptioned image] ) to parameter-efficiently fine-tune the LLM (i.e., Llama [33]), which is responsible to generate the transcriptions given the audio-visual tokens and textual prompt.
 These approaches either employ a single global LoRA to learn audio-visual feature tokens at multiple scales (Multi-Scale LoRA [Uncaptioned image] ), or define multiple LoRAs, each of them focusing on scale-specific audio-visual information (Scale-Specific LoRA [Uncaptioned image] ), or a combination of both (Multi-Scale-Specific LoRA [Uncaptioned image] ).
@@
 Overall, Llama-MTSK exhibits strong performance results, elastic inference, and computational efficiency under a single set of weights.
-Our key contributions are as follows:
-  • We propose Llama-MTSK, the first Matryoshka-based Multimodal LLM designed for audio-visual speech recognition.
-By processing audio-visual tokens with multiple compression levels and granularities, and introducing three Matryoshka-based LoRA modules to efficiently fine-tune the pre-trained LLM, Llama-MTSK is able to dynamically adjust the number of tokens processed during inference using a single model, adapting to varying computational resources or desired accuracy levels.
-  • Llama-MTSK achieves state-of-the-art results on LRS2 and LRS3, the two largest AVSR datasets, consistently exceeding the performance of models independently trained at specific compression levels.
-This trend is observed for the ASR, VSR, and AVSR tasks, across both the evaluated compression techniques and granularities.
 Figure 1: Training and inference stages for Llama-MTSK.
@@
 Refer to caption and Refer to caption represent whether the parameters are trained or kept frozen, respectively.
-2 Llama-MTSK
+Our key contributions are as follows:
+  • We propose Llama-MTSK, the first Matryoshka-based Multimodal LLM designed for audio-visual speech recognition.
+By processing audio-visual tokens with multiple compression levels and granularities, and introducing three Matryoshka-based LoRA modules to efficiently fine-tune the pre-trained LLM, Llama-MTSK is able to dynamically adjust the number of tokens processed during inference using a single model, adapting to varying computational resources or desired accuracy levels.
+  • Llama-MTSK achieves state-of-the-art results on LRS2 and LRS3, the two largest AVSR datasets, consistently exceeding the performance of models independently trained at specific compression levels.
+This trend is observed for the ASR, VSR, and AVSR tasks, across both the evaluated compression techniques and granularities.
+II Llama-MTSK
 The objective of Llama-MTSK is to train an LLM (Llama-based in our setting) that captures audio and visual information at multiple scales, from coarse to fine, thus providing control over the audio-visual granularity during inference.
-Consequently, a single “universal” model allows us to dynamically adjust the performance-efficiency trade-off at inference time, according to specific needs (Cai et al., 2024; Hu et al., 2024a).
-Llama-MTSK follows the structure of Llama-AVSR (Cappellazzo et al., 2025), the first Multimodal LLM (MLLM) tailored for audio-visual speech recognition, with ad-hoc modifications to support MRL (Kusupati et al., 2022).
+Consequently, a single “universal” model allows us to dynamically adjust the performance-efficiency trade-off at inference time, according to specific needs [31, 32].
+Llama-MTSK follows the structure of Llama-AVSR [26], the first Multimodal LLM (MLLM) tailored for audio-visual speech recognition, with ad-hoc modifications to support MRL [28].
 Llama-MTSK computes audio and video tokens via modality-specific pre-trained encoders, and then input them as prefix tokens to the LLM (together with the textual tokens).
-This approach, denoted as decoder-only, is adopted by several architectures due to its versatility and flexibility (Liu et al., 2023; Lin et al., 2024; Fang et al., 2024b; Fan et al., 2024; Zong et al., 2024; Zhang et al., 2025b; Lee et al., 2024; Fini et al., 2024; Li et al., 2024; Tong et al., 2024; Yao et al., 2024).
+This approach, denoted as decoder-only, is adopted by several architectures due to its versatility and flexibility [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
 Llama-MTSK consists of three main components: 1) pre-trained audio and video encoders, 2) audio and video compression and projection modules, and 3) an LLM which is parameter-efficiently fine-tuned via ad-hoc LoRA-based strategies (i.e., LoRA [Uncaptioned image] ).
-2.1 Audio/Video Pre-Trained Encoders
+II-A Audio/Video Pre-Trained Encoders
 We use pre-trained audio and video encoders to project the input audio and video data into two sets of audio and video tokens.
-We denote with $\mathbf{X}^{\mathsf{A}}\in\mathbb{R}^{N_{\mathsf{A}}\times d_{\mathsf{A}}}$ and $\mathbf{X}^{\mathsf{V}}\in\mathbb{R}^{N_{\mathsf{V}}\times d_{\mathsf{V}}}$ the audio and video token sequences, respectively, where $N_{\mathsf{A}}$/$N_{\mathsf{V}}$ is the number of audio/video tokens, and $d_{\mathsf{A}}$/$d_{\mathsf{V}}$ is the audio/video token dimension.
+We denote with $\mathbf{X}^{\mathsf{A}}\in\mathbb{R}^{N_{\mathsf{A}}\times d_{\mathsf{A}}}$ and $\mathbf{X}^{\mathsf{V}}\in\mathbb{R}^{N_{\mathsf{V}}\times d_{\mathsf{V}}}$ the audio and video token sequences, respectively, where $N_{\mathsf{A}}$$N_{\mathsf{V}}$ is the number of audio/video tokens, and $d_{\mathsf{A}}$$d_{\mathsf{V}}$ is the audio/video token dimension.
 The pre-trained encoders are maintained frozen during the training stage ( [Uncaptioned image] in Figure 1).
-2.2 Audio-Visual Compression and Projection
+II-B Audio-Visual Compression and Projection
 Since the dimensions of audio and video tokens often differ from that of the textual tokens, MLLMs include a projection layer that maps audio and video tokens into the LLM embedding space.
-It is common to employ either linear projectors (Liu et al., 2023; Luo et al., 2024; Yao et al., 2024; Li et al., 2024; Liu et al., 2024b; Zhang et al., 2025a) or abstractors (e.g., Q-Former, resampler) (Zhu et al., 2023; Li et al., 2023; Cha et al., 2024).
-In our setting, following (Cappellazzo et al., 2025), we use a two-layer MLP projector.
+It is common to employ either linear projectors [34, 45, 44, 42, 46, 47] or abstractors (e.g., Q-Former, resampler) [48, 49, 50].
+In our setting, following [26], we use a two-layer MLP projector.
 In addition to this, since the LLM predominantly accounts for the entire computation and memory consumption of the MLLM, it is customary to compress the number of multimodal tokens (in our case audio-visual tokens) by a specific factor in order to find the optimal balance in terms of efficiency and accuracy.
-For example, (Cappellazzo et al., 2025; Fang et al., 2024a; Ma et al., 2024; Fathullah et al., 2024) stack multiple consecutive tokens along the token hidden dimension to reduce the number of tokens, whereas other methods rely on the Q-Former architecture (Li et al., 2023) using a fixed number of query tokens (Tang et al., 2023; Yu et al., 2024; Zhang et al., 2025b; Cha et al., 2024).
+For example, [26, 22, 19, 21] stack multiple consecutive tokens along the token hidden dimension to reduce the number of tokens, whereas other methods rely on the Q-Former architecture [49] using a fixed number of query tokens [51, 20, 39, 50].
 However, all these methods need to decide the compression rate to apply beforehand, which means they generate outputs of a single, predetermined length, lacking the ability to modulate the final sequence length.
@@
 With $a_{i}$ we refer both to the compression rate and the corresponding scale interchangeably (e.g., if $a_{i}$ = $4$, then the corresponding sequence would have $\lfloor\frac{N_{\mathsf{A}}}{4}\rfloor$ tokens).
-We then compress the audio and video tokens using the chosen rates, producing token sequences at multiple scales: [$\mathbf{X}^{\mathsf{A}}_{a_{1}},\mathbf{X}^{\mathsf{A}}_{a_{2}},\cdots,\mathbf{X}^{\mathsf{A}}_{a_{\texttt{G}}}$] and [$\mathbf{X}^{\mathsf{V}}_{v_{1}},\mathbf{X}^{\mathsf{V}}_{v_{2}},\cdots,\mathbf{X}^{\mathsf{V}}_{v_{\texttt{T}}}$].
+We then compress the audio and video tokens using the chosen rates, producing token sequences at multiple scales: $\mathbf{X}^{\mathsf{A}}_{a_{1}},\mathbf{X}^{\mathsf{A}}_{a_{2}},\cdots,\mathbf{X}^{\mathsf{A}}_{a_{\texttt{G}}}$] and $\mathbf{X}^{\mathsf{V}}_{v_{1}},\mathbf{X}^{\mathsf{V}}_{v_{2}},\cdots,\mathbf{X}^{\mathsf{V}}_{v_{\texttt{T}}}$].
 At this point, each of these sequences are processed by compression rate-specific linear projectors to align the audio-visual and text tokens (see Figure 1).
-2.3 LLM Adaptation via LoRA [Uncaptioned image]
+II-C LLM Adaptation via LoRA [Uncaptioned image]
 The LLM is responsible for generating the corresponding ASR transcription in an auto-regressive fashion given the audio, video, and textual tokens.
 We define $\mathbf{X}^{\mathsf{AV}}_{ij}$ as the concatenation of audio and video tokens with audio and video compression rates of $a_{i}$ and $v_{j}$, and the prompt textual tokens $\mathbf{X}^{P}$: $\mathbf{X}^{\mathsf{AV}}_{ij}=[\mathbf{X}^{\mathsf{A}}_{a_{i}},\mathbf{X}^{\mathsf{V}}_{v_{j}},\mathbf{X}^{P}]$.
-To parameter-efficiently align the LLM with the multimodal inputs, we use LoRA modules (Hu et al., 2021) to adapt the query and value projection matrices of each layer.
+To parameter-efficiently align the LLM with the multimodal inputs, we use LoRA modules [52] to adapt the query and value projection matrices of each layer.
 In our setting, the LLM is trained on multiple audio-visual tokens with different scales.
@@
 where $s$ is a tunable scalar hyperparameter, $W_{MS}=W_{\texttt{MS}}^{down}W_{\texttt{MS}}^{up}$, $W_{\texttt{MS}}^{down}\in\mathbb{R}^{d\times r}$ and $W_{\texttt{MS}}^{up}\in\mathbb{R}^{r\times d}$, and $r\ll d$ ($r$ is the bottleneck dimension).
-[…]
+In contraposition to MS LoRA [Uncaptioned image] , we propose to learn “expert” LoRA modules, which specialize to each scale.
+We call this approach Specific-Scale (SS) LoRA [Uncaptioned image] .
+Therefore, we define $\texttt{G}\cdot\texttt{T}$ LoRA modules, one for each audio-visual scale.
+We compute the projection output as follows:
 $\mathbf{H}^{\mathsf{AV}}_{ij}\leftarrow\mathbf{X}^{\mathsf{AV}}_{ij}W+s\cdot\mathbf{X}^{\mathsf{AV}}_{ij}W_{\texttt{SS}}^{ij},$ (2)
 where $W_{\texttt{SS}}^{ij}$ is the LoRA decomposition matrix defined for the $i$-th audio scale and $j$-th video scale, and it is defined as $W_{MS}$.
-As we explain in subsection 2.4, while all the LoRA modules are used during the training stage, at inference we only activate one LoRA module, corresponding to the selected audio and video scales.
+As we explain in subsection II-D, while all the LoRA modules are used during the training stage, at inference we only activate one LoRA module, corresponding to the selected audio and video scales.
 The third approach, MSS LoRA [Uncaptioned image] , is a hybrid approach between MS and SS, which aims to learn both scale-specific and multi-scale audio-visual representations.
@@
 Regardless of the LoRA [Uncaptioned image] fine-tuning approach we employ, Llama-MTSK is trained by averaging the auto-regressive next token prediction loss for each audio-visual scale $ij$ for each input data.
-The LLM predicts the response $\mathbf{Y}=\{y_{l}\}_{l=1}^{L}$ conditioned on the multimodal input tokens, where $L$ represents the number of tokens of the ground truth transcription to be generated.
+The LLM predicts the response $\mathbf{Y}=\{y_{l}\}_{l=1}^{L}$ conditioned on the multimodal input tokens, where $L$ is the number of tokens of the ground truth transcription to generate.
 Accordingly, for each Matryoshka audio-visual representation $\mathbf{X}^{\mathsf{AV}}_{ij}$, the probability of the target $\mathbf{Y}$ is computed by:
 $p(\mathbf{Y}|\mathbf{X}^{\mathsf{AV}}_{ij})=\prod_{l=1}^{L}p_{\theta}(y_{l}|\mathbf{X}^{\mathsf{AV}}_{ij},y_{<l}),$ (4)
-where $y_{<l}$ is the generated output sequence up to token $l-1$, and $\theta$ is the trainable parameters, which comprises the projection layers and the LoRA [Uncaptioned image] modules according to the LoRA [Uncaptioned image] fine-tuning approach used.
+where $y_{<l}$ is the generated output sequence up to token l1l-1, and $\theta$ is the trainable parameters, which comprises the projection layers and the LoRA [Uncaptioned image] modules according to the LoRA [Uncaptioned image] fine-tuning approach used.
 The final objective is the average over all the audio-visual token scales:
 $\frac{1}{\texttt{G}\cdot\texttt{T}}\sum_{i=1}^{\texttt{G}}\sum_{j=1}^{\texttt{T}}-\log p(\mathbf{Y}|\mathbf{X}^{\mathsf{AV}}_{ij}).$ (5)
-2.4 Llama-MTSK: Training vs Inference
+II-D Llama-MTSK: Training vs Inference
 During training, Llama-MTSK learns multiple sets of audio-visual tokens, each progressively incorporating more details as the scale increases.
@@
 This is equivalent to one single Llama-AVSR model trained on the specific scale.
-This principle is similar to the behaviour of Mixture of Experts-based models (Shazeer et al., 2017; Fedus et al., 2022; Zoph et al., 2022; Mustafa et al., 2022; Puigcerver et al., 2023; Cappellazzo et al., 2024a; Jiang et al., 2024; Muennighoff et al., 2024), which at inference time only activate a small subset of the available experts (in our case the “experts” are the projectors and LoRA [Uncaptioned image] modules).
+This principle is similar to the behaviour of Mixture of Experts-based models [53, 54, 55, 56, 57, 58, 59, 60], which at inference time only activate a small subset of the available experts (in our case the “experts” are the projectors and LoRA [Uncaptioned image] modules).
 Figure 1 depicts a schematic comparison of Llama-MTSK training and inference processes.
-3 Experiments and Results
-3.1 Implementation Details
+III Experiments and Results
+III-A Implementation Details
 Datasets.
-We train and evaluate Llama-MTSK on LRS2 (Son Chung et al., 2017) and LRS3 (Afouras et al., 2018b), the two largest publicly available datasets for audio-visual speech recognition.
+We train and evaluate Llama-MTSK on LRS2 [61] and LRS3 [62], the two largest publicly available datasets for audio-visual speech recognition.
 LRS2 includes $225$ hours of video clips from BBC programs.
@@
 Pre-Processing.
-We follow (Ma et al., 2023b; Cappellazzo et al., 2025) for the pre-processing of the datasets.
-For the video modality, we crop the mouth region of interests (ROIs) through a bounding box of $96$ $96$.
+We follow [63, 26] for the pre-processing of the datasets.
+For the video modality, we crop the mouth region of interests (ROIs) through a bounding box of $96$ × $96$.
 Each frame is normalised by subtracting the mean and dividing by the standard deviation of the training set.
@@
 Llama-MTSK Details.
-We use Whisper Small and Medium (Radford et al., 2023) as pre-trained audio encoder, whilst AV-HuBERT Large (Shi et al., 2022) for computing the video tokens.
+We use Whisper Small and Medium [64] as pre-trained audio encoder, whilst AV-HuBERT Large [9] for computing the video tokens.
 Their weights remain frozen throughout the training phase.
 The projectors consist of two linear layers with ReLU activation in between.
-As for the LLM, based on the task and dataset, we experiment with $3$ base pre-trained models of varying size from the Llama 3 family (Dubey et al., 2024): Llama 3.1-8B, Llama 3.2-3B, and Llama 3.2-1B.
-Each LoRA module used to fine-tune the query and key projection matrices of each LLM’s self-attention layer has a bottleneck dimension $r$ such that the original LLM’s hidden size is reduced of a factor $32$ for Llama 3.2-3B and 3.2-1B, and $64$ for Llama 3.1-8B (e.g., for Llama 3.2-1B, since the hidden size is $2048$, the rank is set to $2048/32=64$).
+As for the LLM, based on the task and dataset, we experiment with $3$ base pre-trained models of varying size from the Llama 3 family [33]: Llama 3.1-8B, Llama 3.2-3B, and Llama 3.2-1B.
+Each LoRA module used to fine-tune the query and key projection matrices of each LLM’s self-attention layer has a bottleneck dimension $r$ such that the original LLM’s hidden size is reduced of a factor $32$ for Llama 3.2-3B and 3.2-1B, and $64$ for Llama 3.1-8B (e.g., for Llama 3.2-1B, since the hidden size is $2048$, the rank is $2048/32=64$).
 The hyperparameter $s$ is set to $\frac{1}{8}$.
-Table 1: Comparison between Llama-AVSR and our proposed Llama [Uncaptioned image] MS, SS, and MSS approaches on LRS2 and LRS3 benchmarks.
-Llama-AVSR trains $4$ independent models tailored to each configuration of audio-video compression rates.
+TABLE I: Comparison between Llama-AVSR and our proposed Llama [Uncaptioned image] MS, SS, and MSS approaches on LRS2 and LRS3 benchmarks.
+Llama-AVSR trains 44 independent models tailored to each configuration of audio-video compression rates.
 Method Compression Rates (A,V)
@@
 Llama-AVSR 2.4 2.8 3.3 4.1
-[…]
-Llama [Uncaptioned image] SS 2.3 2.2 3.3 3.6
+Llama [Uncaptioned image] MS 2.6 2.7 3.7 4.1
+Llama [Uncaptioned image] $S$ 2.3 2.2 3.3 3.6
 Llama [Uncaptioned image] MSS 2.4 2.4 3.2 3.5
@@
 Llama-AVSR 4.1 4.5 5.3 8.1
-[…]
-Llama [Uncaptioned image] SS 3.4 4.7 4.8 6.4
+Llama [Uncaptioned image] MS 4.8 5.9 6.4 8.9
+Llama [Uncaptioned image] $S$ 3.4 4.7 4.8 6.4
 Llama [Uncaptioned image] MSS 3.6 4.8 6.1 9.0
@@
 We choose the audio and video compression rates to train and evaluate Llama-MTSK carefully, based on the studied tasks.
-For ASR, we apply compression rates in the range of {$4$, $8$, $12$, $16$, $20$}.
-For VSR, since the task is more challenging, we can afford smaller rates: {$1$, $2$, $3$, $4$, $5$} (we also include the case in which no compression is applied).
-For AVSR, we apply audio rates in {$4$, $16$} and video rates in {$2$, $5$}, leading to $4$ audio-visual configurations.
+For ASR, we apply compression rates in the range of {$4$, 88, $12$, $16$, 2020}.
+For VSR, since the task is more challenging, we can afford smaller rates: {$1$, 22, 33, 44, 55} (we also include the case in which no compression is applied).
+For AVSR, we apply audio rates in {$4$, 1616} and video rates in {$2$, 55}, leading to $4$ audio-visual configurations.
 To compress the audio and video tokens, either we apply average pooling with kernel size and stride equal to the desired compression rate, or we stack consecutive frames along the hidden dimension according to the rate (we denote this as “stacking”).
-Table 2: Comparison between Llama [Uncaptioned image] and multiple SOTA methods on the LRS2 and LRS3 benchmarks.
+TABLE II: Comparison between Llama [Uncaptioned image] and multiple SOTA methods on the LRS2 and LRS3 benchmarks.
 The “Lab.
@@
 USR (1,1) 1982/1759 1.9 1.1
-[…]
-[…]
-Llama [Uncaptioned image] SS (4,2) 223/433 2.4 0.9
+Llama-AVSR (4,2) 223/433 2.4 0.9
+Llama [Uncaptioned image] MS (4,2) 223/433 2.1 1.0
+Llama [Uncaptioned image] $S$ (4,2) 223/433 2.4 0.9
 Llama [Uncaptioned image] MSS (4,2) 223/433 2.4 1.2
-[…]
-3.2 AVSR Main Results
-Figure 3: WER results for the average pooling (left) and stacking (right) compression methods for the ASR task. We use Whisper Small as audio encoder and Llama 3.2-1B as LLM.
-We report the results achieved by Llama-MTSK MS, SS, and MSS on the LRS2 and LRS3 datasets in Table 1.
+Figure 3: WER results for the average pooling (left) and stacking (right) compression methods for the ASR task.
+Training/Inference Details.
+Following [26, 63], we augment visual inputs through horizontal flipping, random cropping, and adaptive time masking, while for audio we only apply adaptive time masking.
+For training, we sample babble noise from the NOISEX dataset [65] using a uniform distribution.
+We define the textual prompts as in [26]: “Transcribe {task_prompt} to text.”, where task_prompt $\in$ {“speech”, “video”, “speech and video”}.
+We train our model for $10$ epochs with the AdamW optimizer with cosine annealing scheduler and weight decay set to $0.1$ using NVIDIA A40 GPUs.
+The learning rate is 1e-3 for ASR and AVSR tasks, and 5e-4 for VSR.
+For decoding, we use beam search with a beam width of $15$ and temperature of $0.6$.
+The evaluation metric for all the experiments is the Word Error Rate (WER, %).
+III-B AVSR Main Results
+We report the results achieved by Llama-MTSK MS, SS, and MSS on the LRS2 and LRS3 datasets in Table I.
 We replace “MTSK” with [Uncaptioned image] in the tables and in the following sections to simplify the notation.
@@
 We compare these results with those achieved by training Llama-AVSR independently on the $4$ configurations, leading to $4$ models.
-During inference, Llama-AVSR employs a separate model trained for each audio-video compression rate.
-In contrast, our Llama [Uncaptioned image] uses a single pre-trained model, activating the projector and LoRA [Uncaptioned image] modules corresponding to the desired compression rate.
+During inference, Llama-AVSR employs a separate model trained for each audio-video rate.
+In contrast, our Llama [Uncaptioned image] uses a single pre-trained model, activating the projector and LoRA [Uncaptioned image] modules corresponding to the desired rate.
 On the LRS3 dataset, the three proposed Llama [Uncaptioned image] approaches achieve comparable or superior performance to Llama-AVSR, particularly for the SS and MSS configurations.
-These two methods use LoRA modules specialized for specific compression rates, which are activated during inference based on specific requirements.
-On the LRS2 dataset, Llama [Uncaptioned image] SS outperforms all other approaches across all compression rates.
+These two methods use LoRA modules specialized for specific rates, which are activated during inference based on specific requirements.
+On the LRS2 dataset, Llama [Uncaptioned image] SS outperforms all other approaches across all rates.
 Llama [Uncaptioned image] vs SOTA Methods.
-In Table 2, we compare Llama [Uncaptioned image] with state-of-the-art (SOTA) methods on LRS2 and LRS3 for the AVSR task.
+In Table II, we compare Llama [Uncaptioned image] with state-of-the-art (SOTA) methods on LRS2 and LRS3 for the AVSR task.
 We equip Llama [Uncaptioned image] with Whisper Medium and Llama 3.1-8B.
-We report results from $5$ recent SOTA AVSR methods: CM-seq2seq (Ma et al., 2021), Efficient Conformer (Burchi and Timofte, 2023), auto-avsr (Ma et al., 2023b), Whisper-Flamingo (Rouditchenko et al., 2024), and USR (Haliassos et al., 2024a).
+We report results from $5$ recent SOTA AVSR methods: CM-seq2seq [5], Efficient Conformer [66], auto-avsr [63], Whisper-Flamingo [67], and USR [13].
 Notably, all these methods do not reduce the token sequence length, whereas Llama-AVSR and Llama [Uncaptioned image] reduce the number of tokens by a factor $4$ for audio and $2$ for video.
@@
 Finally, more AVSR experiments can be found in the Appendix.
-3.3 Additional Results
+III-C Additional Results
 In this section, we extend our analysis to the tasks of ASR and VSR, where only audio or video tokens are fed to the LLM, respectively.
 We finally present the computational cost analysis of Llama [Uncaptioned image] .
-Table 3: Comparison between Llama [Uncaptioned image] MS and a training-free Llama-AVSR-based approach that reduces the number of tokens via average pooling at inference time for the ASR task on the LRS3 dataset.
+Figure 4: WER results for the average pooling (left) and stacking (right) compression methods for the VSR task. We use AVHuBERT Large as video encoder and Llama 3.2-3B as LLM.
+ASR Results.
+For the ASR task, we consider $5$ compression rates in the range {$4$, 88, $12$, $16$, 2020}.
+In Figure 3, we report the results on the LRS3 dataset when using average pooling compression (left) and stacking compression (right).
+With the exception of rate = $20$, all the three Llama [Uncaptioned image] methods outperform separately-trained Llama-AVSR methods.
+The MSS configuration achieves the best WER performance across all the compression rates, even surpassing or equaling the performance of Llama-AVSR trained at the lowest compression rate of $20$.
+TABLE III: Comparison between Llama [Uncaptioned image] MS and a training-free Llama-AVSR-based approach that reduces the tokens via average pooling at inference time for the ASR task on LRS3.
 Method Compression Rate
@@
 Avg Pooling 4.3 13.5 46.1 89.2 160.0
-[…]
-Figure 4: WER results for the average pooling (left) and stacking (right) compression methods for the VSR task. We use AVHuBERT Large as video encoder and Llama 3.2-3B as LLM.
-ASR Results.
-For the ASR task, we consider $5$ compression rates in the range {$4$, $8$, $12$, $16$, $20$}.
-In Figure 3, we report the results on the LRS3 dataset when using average pooling compression (left) and stacking compression (right).
-With the exception of rate = $20$, all the three Llama [Uncaptioned image] methods outperform separately-trained Llama-AVSR methods.
-The MSS configuration achieves the best WER performance across all the compression rates, even surpassing or equaling the performance of Llama-AVSR trained at the lowest compression rate of $20$.
+Llama [Uncaptioned image] MS 2.5 2.3 2.3 2.7 3.0
 VSR Results.
 Figure 4 shows WER results for the VSR task, similar to the ASR results in Figure 3.
-The video rates are {$1$, $2$, $3$, $4$, $5$}, lower than the ASR rates due to the greater complexity of VSR.
+The video rates are {$1$, 22, 33, 44, 55}, lower than the ASR rates due to the greater complexity of VSR.
 For both average pooling and stack compression, all three Llama [Uncaptioned image] approaches outperform Llama-AVSR, with increasing gains at higher rates.
@@
 However, all of them significantly outperform Llama-AVSR.
-Table 4: Computational cost analysis of Llama [Uncaptioned image] MS using different compression rates and Llama 3.1-8B.
-(A,V) Rates # Tokens TFLOPs
-(1,1) 757 11.40
-[…]
-(4,5) 182 2.74
-(16,2) 163 2.46
-(16,5) 88 1.33
 Llama [Uncaptioned image] vs Avg Pooling at Inference Time.
@@
 We compare our method with a training-free approach that trains a single Llama-AVSR model without compression and then applies the desired compression rate at inference on-the-fly by average pooling the tokens.
-In Table 3, we study the ASR setting with audio compression rates in the range {$2$, $4$, $6$, $8$, $10$}.
+In Table III, we study the ASR setting with audio rates in the range {$2$, 44, 66, 88, 1010}.
 The performance of the average-pooling baseline is severely impacted by a decrease in the number of tokens, while Llama [Uncaptioned image] MS is much more robust.
 These results demonstrate that Llama [Uncaptioned image] MS can be effectively used with diverse computational resources.
-Notably, even with limited resources, a compression rate of $8$ incurs minimal performance loss.
 Computation Cost Analysis.
-Table 4 presents the computational benefits of using Llama [Uncaptioned image] .
-Specifically, we evaluate MS LoRA [Uncaptioned image] with Llama 3.1-8B as LLM and detail the associated inference costs.
-Without compression, we assume the LLM processes $500$ audio tokens, $250$ video tokens (the resolution of the audio encoder is twice that of the video encoder), and $7$ tokens for the textual prompt, totaling $757$.
-As shown in the table, our proposed approach yields speedups, reducing TFLOPs by over $8$x when applying compression rates of $16$ and $5$ for audio and video, respectively, thus substantially improving efficiency.
-4 Conclusion
-This work introduces Llama-MTSK, a versatile audio-visual MLLM capable of elastic inference across multiple tasks and computational resources.
-Llama-MTSK exploits the concept of matryoshka representation learning to adapt the pre-trained LLM through ad-hoc LoRA [Uncaptioned image] modules, achieving performance comparable to or better than models separately trained on each compression rate while significantly reducing computational costs.
-5 Limitations
-[…]
+In Figure 5, we illustrate the benefits of Llama [Uncaptioned image] MS in terms of TFLOPS and inference costs.
+Without compression (i.e., (1,1)), we assume the LLM processes $500$ audio tokens, $250$ video tokens (the resolution of the audio encoder is twice that of the video encoder), and $7$ tokens for the textual prompt, totaling $757$.
+By increasing the audio-visual compression rates, we reduce the number of tokens processed by the LLM, and thus the TFLOPs, by up to 8.68.6x when applying compression rates of (16,5), resulting in $88$ tokens.
+Despite this substantial reduction in TFLOPs, the resulting increase in WER remains modest.
+Moreover, Llama [Uncaptioned image] MS enables elastic inference by allowing users to select compression rates based on their computational constraints.
+Figure 5: Comparison of Llama Refer to caption MS in terms of number of audio-visual processed tokens, WER, and TFLOPS.
+IV Conclusion
+We introduce Llama-MTSK, a versatile audio-visual MLLM capable of elastic inference across multiple tasks and computational resources.
+Llama-MTSK exploits matryoshka representation learning to adapt the pre-trained LLM through ad-hoc LoRA [Uncaptioned image] modules, achieving performance comparable to or better than models separately trained on each compression rate while significantly reducing computational costs.
 Appendix A Appendix
-Figure 5: Additional WER results using stacking compression for the ASR task with {$2$, $4$, $6$, $8$, $10$} rates. We use the same configuration as in Figure 3.
-A.1 Additional Experiments for ASR
-In this section, we report additional results for the ASR task when using compression rates in a different range , specifically {$2$, $4$, $6$, $8$, $10$}.
-Compared to Figure 3, the increment between two consecutive rates is halved.
-We argue that it is more useful to use more diverse rates for ASR since we do not observe much deterioration of the WER results when doubling the rate (in Figure 5, the baseline Llama-AVSR achieves similar results when compressing the tokens of a factor $2$, $4$, and $6$).
+Figure 6: Additional WER results using stacking compression for the ASR task with {22, 44, 66, 88, 1010} rates.
+Appendix B Additional Experiments for ASR
+Figure 7: Additional results for Llama Refer to caption using stacking compression on the LRS3 dataset.
+In this section, we report additional results for the ASR task when using compression rates in a different range, specifically {$2$, 44, 66, 88, 1010}.
+Compared to Figure 3 in the main paper, the increment between two consecutive rates is halved.
+We argue that it is more useful to use more diverse rates for ASR since we do not observe much deterioration of the WER results when doubling the rate (in Figure 6, the baseline Llama-AVSR achieves similar results when compressing the tokens of a factor $2$, 44, and $6$).
 Figure 6 shows that Llama [Uncaptioned image] MS and MSS achieves comparable or better performance than Llama-AVSR.
@@
 This argument also explains why for rates $8$ and $10$ the MS variant performs better than the other ones.
-Figure 6: Additional results for Llama Refer to caption using stacking compression on the LRS3 dataset.
-A.2 AVSR Results with Stacking Compression
+Appendix C AVSR Results with Stacking Compression
 We include additional results for AVSR on LRS3 using the stacking compression method in Figure 7.
@@
 Our three proposed Matryoshka approaches performs better than or equally well as Llama-AVSR, especially under conditions of high audio compression, underscoring the effectiveness of our proposed Llama [Uncaptioned image] .
-A.3 Full AVSR Results with Whisper Medium and LLama 3.1-8B
-In Table 2, we only included for Llama-AVSR and Llama [Uncaptioned image] the results with audio and video compression rates equal to $4$ and $2$, respectively.
-In Table 5, we also report additional configurations of audio-video compression rates.
+Appendix D Full AVSR Results with Whisper Medium and LLama 3.1-8B
+In Table 2 in the main paper, we only included for Llama-AVSR and Llama [Uncaptioned image] the results with audio and video compression rates equal to $4$ and $2$, respectively.
+In Table IV, we also report additional configurations of audio-video compression rates.
 We use Whisper medium as audio encoder and Llama 3.1-8B as LLM.
@@
 In particular, we highlight the sizeable gains brought by all the three LoRA [Uncaptioned image] approaches for LRS3 when we apply the highest compression rates configuration (16,5).
-Table 5: Comparison between Llama-AVSR and our proposed Llama [Uncaptioned image] MS, SS, and MSS approaches on LRS2 and LRS3 benchmarks.
+TABLE IV: Comparison between Llama-AVSR and our proposed Llama [Uncaptioned image] MS, SS, and MSS approaches on LRS2 and LRS3 benchmarks.
 We employ Whisper medium and Llama 3.1-8B.
-Llama-AVSR trains $4$ independent models tailored to each configuration of audio-video compression rates.
+Llama-AVSR trains 44 independent models tailored to each configuration of audio-video compression rates.
 Method Compression Rates (A,V)
@@
 Llama-AVSR 0.9 0.9 1.6 2.1
-[…]
-Llama [Uncaptioned image] SS 0.9 1.0 1.7 1.8
+Llama [Uncaptioned image] MS 1.0 1.1 1.5 1.6
+Llama [Uncaptioned image] $S$ 0.9 1.0 1.7 1.8
 Llama [Uncaptioned image] MSS 1.2 1.0 1.5 1.6
@@
 Llama-AVSR 2.4 2.2 2.9 3.3
-[…]
-Llama [Uncaptioned image] SS 2.4 2.1 2.9 2.9
+Llama [Uncaptioned image] MS 2.1 2.3 2.9 3.2
+Llama [Uncaptioned image] $S$ 2.4 2.1 2.9 2.9
 Llama [Uncaptioned image] MSS 2.4 2.5 3.2 3.4